OpenStudy (starryknight):

What is the solution to the following system of equations? (metal will be given to person who helps me) -8x - 6y = 52 8x + 0y = -16

OpenStudy (anonymous):

does it really say \(0y\) ??!!

OpenStudy (anonymous):

must be a computer generated question

OpenStudy (starryknight):

my hw said exactly that

OpenStudy (starryknight):

i take it online so its hard to understand the homework

OpenStudy (mathmale):

Hello, SK: I assume you know that a "solution" here signifies a point, (x, y), whose coordinates satisfy the given set of equations. Our job is to find such a solution. I'm glad satellite has pointed out that that coefficient of " 0 " for y appears to be unusual.

OpenStudy (mathmale):

With such a system, we need to eliminate one variable and solve the resulting equation for the other. In -8x - 6y = 52 8x + 0y = -16 note that -8x+8x = 0. Please combine all like terms: -8x+8x=? -6y+0y=? 52-16= ?

OpenStudy (starryknight):

x^2-6y^2=36 ? did i do it ok :((

OpenStudy (mathmale):

Let's combine: -8x - 6y = 52 8x + 0y = -16 ------------- 0 -6y = 36 Please solve this resulting equation for y.

OpenStudy (starryknight):

y=-6

OpenStudy (mathmale):

Remember: "combine" signifies addition or subtraction. No squares result. Perfect. Now that you have y=-6, how would you go about finding the value of x?

OpenStudy (mathmale):

Hint: you have two equations relating x and y: -8x - 6y = 52 8x + 0y = -16 You've found that y=-6. You could: (1) substitute that y=-6 into the first equation and solve for x, or (2) note that the second equation simplifies to 8x=-16 and solve this for x. y = -6 x = ?

OpenStudy (starryknight):

-2 is x

OpenStudy (starryknight):

?

OpenStudy (mathmale):

Very sorry for my slow response. Yes. x=-2. Now, having obtained x=-2 and y=-6, how would you go about checking your answer?